6-4(-6x+8)=7(-3x-3)+7
Answer by LarissaRichardson(34)

First Distribute
6-4(-6x+8)=7(-3x-3)+7
To get rid of th parenthesis multiply the -4(-6x+8) and 7(-3x-3)+7
To get:
6-24x-32=-21x-21+7
Now combine like terms
-24x-26=-21x-14
And then continue to combine like terms
-26+14=-21x+24x
-12=3x
Divide by 3
-4=x